helpQuestion 11 The classic five-stage pipeline MIPS architecture is used to execute the code fragments in this problem. Assume t
Answer #1

value of R4 = 200
Number of times Loop executes = 200/4 = 50

(a)size of 1 instruction = 4 bytes
size of memory needed = 9*4 = 36 bytes

(b)Number of times Loop executes = 200/4 = 50

(c)Total number of instructions executed = 50*9 = 450

(d)Total number of data memory for read in one iteration = 2

(e)Total number of data memory for write in one iteration = 2

(f)Total number of instruction memory in one iteration = 9